OpenStudy (anonymous):

Compare and contrast the two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1- (4,4) and (2,6) Set #2- (2,4) and (4,2) A) The slope of Set #1 is smaller than the slope of Set #2. B) The slope of Set #1 is larger than the slope of Set #2. C) The slope of Set #1 is the same as the slope of Set #2. D) None of the statements above describes slopes of sets shown.

OpenStudy (texaschic101):

we will find the slope of each using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) set 1 (4,4) x1 = 4 and y1 = 4 (2,6) x2 = 2 and y2 = 6 now we sub slope(m) = (6 - 4) / (2 - 4) m = 2/-2 m = -1 set 2 (2,4) x1 = 2 and y1 = 4 (4,2) x2 = 4 and y2 = 2 now we sub slope(m) = (2 - 4) / (4 - 2) m = -2/2 m = -1 Slope of set 1 and set 2 are the same